OpenStudy (august899):

Factor the polynomials:m^3+1/8 and p^3+216

zepdrix (zepdrix):

Hey :)

zepdrix (zepdrix):

\[\large\rm m^3+\frac{1}{8}\]Hmmm, recall that we have a rule for factoring the `sum of cubes`: \[\large\rm a^3+b^2=(a+b)(a^2-ab+b^2)\]

zepdrix (zepdrix):

But with the way our problem is written right now, we only have one cube. Any ideas how to turn the 1/8 into `something`^3 ?

zepdrix (zepdrix):

@August899

OpenStudy (august899):

do we divide 1 over 8?

zepdrix (zepdrix):

The 8 is some number to the third power.

zepdrix (zepdrix):

Example: 27 is 3 to the third power. Because 3 x 3 x 3 = 27 Therefore \(\large\rm 27=3^3\)

zepdrix (zepdrix):

Comeon Auguuuuuust >.< break down the 8, how can we write that another way?